# Balancing properties of tropical moduli maps

### v1

@@ Abstract.
KC was partially supported by NSF FRG grant DMS–2053261 and a Minerva Short-Term Research Grant. XH is supported by the National Key R&\&&D Program of China 2022YFA1007100, the NSFC grant 12301057, and the ERC Consolidator Grant 770922 - BirNonArchGeom.
@@ 1. Introduction
In this paper, we study certain properties of tropicalizations of families of parameterized curves over a valued field $K$. Given a family of smooth curves $f\colon{\mathcal{C}}\to U$ together with a map ${\mathcal{C}}\to X$ to a toric variety $X$, we construct a family of tropical curves $\Gamma_{\Lambda}\to\Lambda$ together with a map $h\colon\Gamma_{\Lambda}\to{\mathbb{R}}^{n}$. This generalizes the construction for $1$-dimensional families in our previous paper [CHT23] and relies on results of de Jong [dJ96] on the existence of semistable models and of Gubler, Rabinoff, and Werner [GRW16] on skeletons of such models.
While other versions of tropicalizations of families of curves have been established in the literature, see for example [ACGS20, CCUW20, Ran22, CGM22], we propose here to focus on properties of the induced moduli map. Namely, the family $\Gamma_{\Lambda}\to\Lambda$ induces a tropical moduli map
$\alpha\colon\Lambda\to M_{g,n,\nabla}^{\mathrm{trop}}$
to the moduli space $M_{g,n,\nabla}^{\mathrm{trop}}$ of tropical parameterized curves of genus $g$, degree $\nabla$, and with $n$ marked points. As usual, it is given by sending a point in $\Lambda$ to the isomorphism class of its fiber in the tropical family. One could expect such a tropical moduli map to be a tropical map itself, i.e., to satisfy a balancing condition. Our main results give first evidence for this expectation.
Before we can state these results, recall that $M_{g,n,\nabla}^{\mathrm{trop}}$ is a (generalized) polyhedral complex with polyhedra $M_{[\Theta]}$ parameterizing tropical curves of fixed combinatorial type $\Theta$. Two classes of combinatorial types are of particular interest for our purposes:
First, weightless and $3$-valent types. The corresponding strata in $M_{g,n,\nabla}^{\mathrm{trop}}$ are maximal with respect to inclusion in the closure. Thus one may think of them as parameterizing tropically general curves.
Second, weightless and almost $3$-valent types, where we allow for a single $4$-valent vertex. A fixed weightless and almost $3$-valent stratum is contained in the closure of at most three other strata: the $3$ ways to resolve the $4$-valent vertex into a pair of adjacent $3$-valent vertices, each giving rise to a potential adjacent stratum some of which might get identified by an automorphism of the combinatorial type. Notice that the slope of the resolving edge is determined uniquely by the harmonicity condition, and each of these strata is weightless and $3$-valent.
Here is an informal version of our main result; see Theorem 3.1 for a precise statement:
@@
Let $\alpha\colon\Lambda\to M_{g,n,\nabla}^{\mathrm{trop}}$ be as above induced by the tropicalization of a family of parameterized curves. Then,
@@
    The map $\alpha$ is (quasi-)harmonic on the interior of weightless and 3-valent strata.
@@
    The map $\alpha$ is either (quasi-)harmonic or locally combinatorially surjective on the interior of weightless and almost 3-valent strata.
More precisely, the map $\alpha$ is harmonic on codimension $1$ strata of $\Lambda$ and quasi-harmonic on strata of higher codimension. Here, quasi-harmonic means that a positive linear combination of the images of the primitive normal vectors to a stratum is contained in the image of the stratum; harmonic requires that the coefficients in the linear combination can be chosen to be $1$. Locally combinatorially surjective, on the other hand, means that the image of $\alpha$ contains points in each of the weightless and 3-valent strata adjacent to the weightless and almost 3-valent stratum. See Definitions 2.4 and 2.9.
Remark 1.1.
In [CHT22] we also considered types $\Theta$ that are not weightless or almost $3$-valent. Namely, ones that contain either a single vertex of weight $1$ or two vertices of valence $4$ adjacent to a flattened cycle. We established the corresponding liftability results for such strata in [CHT22, Lemmas 4.5 and 4.6] in the case of a $1$-dimensional family of parameterized curves. Using the arguments we present in the current article, one can also generalize these results to the setting of higher dimensional bases.
The Main Theorem allows us, in particular, to establish the following corollary, which we again state informally and refer to Corollary 3.2 for a precise statement:
@@
With assumptions as in the Main Theorem we have:
@@
    If $\alpha$ is locally surjective at a point in the interior of a weightless and 3-valent stratum, it is surjective onto the closure of that stratum.
@@
    If $M_{[\Theta]}$ is a weightless and almost 3-valent stratum and $\alpha$ is surjective onto one of its adjacent weightless and 3-valent strata, it is surjective also onto all other such adjacent strata.
We view this result as a new kind of liftability/realizability criterion for parameterized tropical curves. To illustrate this, let us explain how we will use the corollary in [CHT24]. In loc. cit. we apply the corollary to the universal family ${\mathcal{C}}\to V$ over an irreducible component $V$ of a Severi variety of a toric surface $X$. Given a general point $[C]\in V$, parameterizing a curve of geometric genus $g$ on $X$, we can cut out an $N$-dimensional neighborhood of $[C]$ in $V$ by varying $N$ general point constraints, where $N$ is the dimension of $V$. Choosing the points general enough ensures that the combinatorial type $\Theta$ of the tropicalization $\mathrm{trop}(C)$ of $C$ is weightless and 3-valent, and the tropicalization of the $N$ points cuts out an $N$-dimensional neighborhood of $\mathrm{trop}(C)$ in $M_{[\Theta]}$. As it turns out, $\dim M_{[\Theta]}=N$, and hence part (1) of the corollary applies. Using part (2), it follows that the tropicalization of the universal family over $V$ contains curves in the closure of any weightless and 3-valent stratum that can be reached from $M_{[\Theta]}$ by passing through weightless and (almost) 3-valent strata.
Thus, the question of whether a given tropical curve is the tropicalization of a curve in ${\mathcal{C}}\to V$ is reduced to a purely combinatorial question, reminiscent of establishing connectivity in codimension $1$. Consequently, we show for a large class of toric surfaces that each component $V$ contains a point representing a curve with a certain tropicalization. In a second step, we show by deformation-theoretic arguments that curves with such a tropicalization can be contained in at most one irreducible component $V$, which allows us to conclude the irreducibility of Severi varieties for these toric surfaces.
@@ 2. Notation and terminology
Throughout this paper, $F$ is a complete discretely valued field with an algebraically closed residue field, and $K$ is its algebraic closure. We denote the valuation by $\nu\colon K\rightarrow\mathbb{R}\cup\{\infty\}$, the ring of integers by $K^{0}$, its maximal ideal by $K^{00}$, and the residue field by $\widetilde{K}$, and similarly for $F$.
For a subset $P\subset{\mathbb{R}}^{n}$ we denote by ${\rm Lin}(P)$ the linear subspace acting simply transitively on the affine hull of $P$. If $P$ is a polyhedron with rational slopes, then ${\rm Lin}(P)$ admits a natural integral structure induced from ${\mathbb{R}}^{n}$, i.e., the lattice of integral vectors. By an integral structure on $P$, we mean the integral structure on ${\rm Lin}(P)$.
@@ 2.1. Polyhedral complexes
An (abstract) polyhedral complex with integral structure is a connected topological space $\Lambda$ together with a finite set $\mathcal{F}(\Lambda)$ of closed subsets, called faces of $\Lambda$, and a map $f_{W}\colon W\rightarrow N_{W}\otimes{\mathbb{R}}$ for every $W\in\mathcal{F}(\Lambda)$ such that
@@
    For every $W\in\mathcal{F}(\Lambda)$, $N_{W}$ is a lattice of finite rank;
@@
    The map $f_{W}$ is a homeomorphism from $W$ to a polyhedron with rational slopes of dimension ${\rm rank}(N_{W})$. By abuse of language, we will address $W$ itself as a polyhedron and speak about its faces, interior, dimension, integral structure, etc.;
@@
    If $W\in\mathcal{F}(\Lambda)$ and $W^{\prime}\in\mathcal{F}(\Lambda)$ is a face of $W$, then the integral structure on $W^{\prime}$ is the restriction of the integral structure on $W$, i.e., the linear part of $f_{W}\circ f_{W^{\prime}}^{-1}$ takes $N_{W^{\prime}}$ to a saturated sublattice in $N_{W}$.
We say that $\Lambda$ is rational if each $f_{W}(W)$ has rational vertices in $N_{W}\otimes{\mathbb{R}}$. All polyhedral complexes considered in this paper will be assumed to have integral structures, and we will call them polyhedral complexes for short.
@@
A map $f\colon\Lambda\rightarrow\Lambda^{\prime}$ is called piecewise integral affine if for each face $W\in\mathcal{F}(\Lambda)$ there is a face $W^{\prime}\in\mathcal{F}(\Lambda^{\prime})$ such that $f(W)\subseteq W^{\prime}$ and $f|_{W}\colon W\to W^{\prime}$ is integral affine.
For a face $W\in\mathcal{F}(\Lambda)$, let $\{W_{i}\}_{i}$ be the set of faces in $\mathcal{F}(\Lambda)$ containing $W$ as a codimension-one face. Then, for each $i$, the lattice $N_{W}$ is naturally identified with a saturated sublattice of $N_{W_{i}}$ of corank one. Let $\vec{e}_{i}+N_{W}$ be the generator of $N_{W_{i}}/N_{W}$ for which $f_{W_{i}}(W_{i})\subseteq f_{W_{i}}(W)+\mathbb{R}_{\geq 0}\vec{e}_{i}$, and denote by ${\rm Star}(W)$ the collection $\{\vec{e}_{i}+N_{W}\}$. We will often consider a small open neighborhood of $W^{\circ}$ in $W^{\circ}\cup\bigcup_{i}W_{i}^{\circ}$. To distinguish this set from ${\rm Star}(W)$, we denote it by ${\rm Star}(W^{\circ})$.
@@
Let $\Lambda$ be a polyhedral complex and $W\in\mathcal{F}(\Lambda)$ a face. A piecewise integral affine map $f\colon\Lambda\rightarrow{\mathbb{R}}^{n}$ is called quasi-harmonic at $W$ if
$\sum_{\vec{e}+N_{W}\in{\rm Star}(W)}a_{\vec{e}}\cdot\frac{\partial f}{\partial\vec{e}}\in{\rm Lin}(f(W))$
for some positive integers $a_{\vec{e}}$. It is called harmonic at $W$ if one can choose all $a_{\vec{e}}=1$. Finally, the map $f$ is called (quasi-)harmonic if it is (quasi-)harmonic at all codimension-one faces.
@@
Plainly, being (quasi-)harmonic at $W$ depends only on the restriction of the map $f$ to ${\rm Star}(W^{\circ})$.
@@ 2.2.1. Abstract and parameterized tropical curves
The graphs in this paper are finite graphs having vertices, edges, and half-edges, called legs. We denote the set of vertices of a graph $\mathbb{G}$ by $V({\mathbb{G}})$, edges by $E({\mathbb{G}})$, and legs by $L({\mathbb{G}})$. We set $\overline{E}(\mathbb{G}):=E({\mathbb{G}})\cup L({\mathbb{G}})$. By a tropical curve $\Gamma$ we mean a finite graph $\mathbb{G}$ with ordered legs equipped with a length function $\ell\colon E(\mathbb{G})\rightarrow\mathbb{R}_{>0}$ and a weight (or genus) function $g\colon V(\mathbb{G})\rightarrow\mathbb{Z}_{\geq 0}$. We say that a curve is weightless if the weight function is identically zero. For any leg $l\in L({\mathbb{G}})$ we set $\ell(l):=\infty$.
For $e\in\overline{E}({\mathbb{G}})$, we use $\vec{e}$ to indicate a choice of orientation on $e$. If $e\in L({\mathbb{G}})$ is a leg, then it will always be oriented away from the vertex. Bounded edges will be considered with both possible orientations. We can always view tropical curves as polyhedral complexes by dividing the loops with the middle point, identifying the edges of ${\mathbb{G}}$ with bounded closed intervals of the corresponding lengths, and identifying the legs of ${\mathbb{G}}$ with semi-bounded closed intervals. This way, ${\rm Star}(v)$ is in one-to-one correspondence with the collection of oriented edges and legs having $v$ as their tail. In particular, ${\rm Star}(v)$ contains two vectors/edges for every loop based at $v$. The number of elements in ${\rm Star}(v)$ is called the valence of $v$. The genus of $\Gamma$ is defined to be $g(\Gamma)=g({\mathbb{G}}):=1-\chi(\mathbb{G})+\sum_{v\in V(\mathbb{G})}g(v)$, where $\chi({\mathbb{G}}):=b_{0}({\mathbb{G}})-b_{1}({\mathbb{G}})$ denotes the Euler characteristic of ${\mathbb{G}}$. Finally, a tropical curve $\Gamma$ is called stable if $|{\rm Star}(v)|+2g(v)\geq 3$ for every vertex $v\in V({\mathbb{G}})$.
Let $N$ be a lattice and set $N_{\mathbb{R}}:=N\otimes{\mathbb{R}}$. In the sequel, we will always let $N$ be the lattice of cocharacters of a toric variety. A parameterized tropical curve is a harmonic piecewise integral affine map $h\colon\Gamma\rightarrow N_{\mathbb{R}}$ from a tropical curve $\Gamma$ (considered as a polyhedral complex) to $N_{\mathbb{R}}$. The combinatorial type $\Theta$ of a parameterized tropical curve is defined to be the weighted underlying graph $\mathbb{G}$ with ordered legs equipped with the collection of slopes $\frac{\partial h}{\partial\vec{e}}$ for $e\in\overline{E}({\mathbb{G}})$. We denote the group of automorphisms of a combinatorial type $\Theta$ by $\mathrm{Aut}(\Theta)$, and the isomorphism class of $\Theta$ by $[\Theta]$. We define the extended degree $\overline{\nabla}$ of a parameterized tropical curve to be the sequence of slopes $\left(\frac{\partial h}{\partial\vec{l}}\right)_{l\in L({\mathbb{G}})}$, and the degree $\nabla$ the subsequence $\left(\frac{\partial h}{\partial\vec{l}_{i}}\right)$ of non-zero slopes.

### v2

@@ Abstract.
KC was partially supported by NSF FRG grant DMS–2053261, a Minerva Short-Term Research Grant, and the Center for Advanced Studies at BGU. XH is supported by the National Key R&\&D Program of China 2022YFA1007100, the NSFC grant 12301057, and the ERC Consolidator Grant 770922 - BirNonArchGeom.
2020 Mathematics Subject Classification: Primary – 14T20; Secondary – 14H10
@@ 1. Introduction
Starting with the seminal work of Mikhalkin, [Mik05], tropical geometry emerged as a very useful tool in many aspects of algebraic geometry, including enumerative problems, questions about the geometry of various moduli spaces, etc. Tropical varieties are polyhedral objects controlling degenerations of algebraic varieties, and tropical methods in algebraic geometry can be considered as a refined version of degeneration and deformation-theoretic techniques.
Most applications of tropical geometry build upon so-called tropicalization and lifting results. Roughly speaking, tropicalization is a kind of linearization procedure associating a tropical object to an algebraic one, and lifting results aim at realizing a given tropical object as the tropicalization of an algebro-geometric one. There exists no functorial way to tropicalize algebraic geometry. Yet, various useful tropicalization constructions have been developed, see, e.g., [Mik05, Tyo12, ACP15, ABBR15, Tyo17, ACGS20, CCUW20, Ran22, MUW21, CGM22, CHT23]. In particular, in [CHT23], we introduced tropicalizations of one-parameter families of curves in toric varieties and studied their properties. As an application, in loc. cit. we resolved the Severi problem in arbitrary characteristic.
In general, lifting results are more challenging to achieve as they are based on deformation theory, which often is obstructed and difficult to control. Surprisingly, the basic balancing properties of tropicalizations of families established in [CHT23] allowed us to prove non-trivial lifting results without any deformation theory. The goal of the current paper is to extend the theory initiated in loc. cit. to families of curves over higher-dimensional bases. In [CHT25, CHT26], we use the results of the current paper to prove the irreducibility of Hurwitz spaces in arbitrary characteristic, and to count moduli of curves of given genus that can be realized as a curve in a given linear system on a toric surface. We believe that the results of the current paper will find further applications in the future.
Let us describe the main results of the paper in more detail. Suppose we are given a family of smooth curves ${\mathcal{C}}\to U$ with marked points together with a map $f\colon{\mathcal{C}}\to X$ to a toric variety $X$. After replacing $U$ with an appropriate alteration, we construct a family of stable curves whose base $B$ admits a natural tropicalization $\Lambda$ in the sense of Gubler, Rabinoff, and Werner [GRW16]. We then construct a family of parameterized tropical curves $\left(\Gamma_{\Lambda}\to\Lambda,h\colon\Gamma_{\Lambda}\to{\mathbb{R}}^{n}\right)$ such that for any $b\in B$ the tropicalization of $f_{b}\colon{\mathcal{C}}_{b}\to X$ is the fiber $h_{q}\colon\Gamma_{q}\to{\mathbb{R}}^{n}$, where $q=\mathrm{trop}(b)\in\Lambda$. The family of parameterized tropical curves $\left(\Gamma_{\Lambda}\to\Lambda,h\colon\Gamma_{\Lambda}\to{\mathbb{R}}^{n}\right)$ induces a natural map to the moduli space of stable parameterized tropical curves of appropriate degree $\nabla$, genus $g$, and number $n$ of contracted legs
$\alpha\colon\Lambda\to M_{g,n,\nabla}^{\mathrm{trop}},$
given by $q\mapsto(\Gamma_{q},h_{q})$; see § 2.2.2 for the description of the moduli space $M_{g,n,\nabla}^{\mathrm{trop}}$ and § 2.2.3 for the definition of families of parameterized tropical curves and the construction of the induced map $\alpha$.
It is natural to expect this map to be itself a tropical map, i.e., a piecewise integral affine map satisfying certain balancing properties. Our main result provides first evidence for this expectation. Namely, we show that the map $\alpha$ is a piecewise integral affine map, and that it satisfies natural balancing properties along two types of strata of small codimension: the maximal strata classifying weightless and $3$-valent parameterized tropical curves, and their codimension-one boundary strata classifying weightless parameterized tropical curves which are almost $3$-valent, i.e., $3$-valent except for a unique $4$-valent vertex. Here is our main result (see Theorem 3.1 for the precise formulation).
@@
[…]
@@
    For any face $W$ of $\Lambda$ of codimension one (resp. of codimension at least two) such that the tropical curves over the interior of $W$ are weightless and $3$-valent except for at most one $4$-valent vertex, the induced map $\alpha\colon\Lambda\to M^{\mathrm{trop}}_{g,n,\nabla}$ is either harmonic (resp. quasi-harmonic) or locally combinatorially surjective along $W$.
Here, quasi-harmonic means that a positive linear combination of the images of the primitive normal vectors to a stratum is contained in the image of the stratum, and harmonic requires that the coefficients in the linear combination can be chosen to be $1$; see Definition 2.4. Locally combinatorially surjective, on the other hand, means that the image of $\alpha$ contains points in each of the weightless and $3$-valent strata adjacent to the weightless and almost $3$-valent stratum; see Definition 2.10.
As a consequence of the Main Theorem we establish the following corollary (see Corollary 3.3) that provides a new relative realizability criterion for (parameterized) tropical curves.
@@
Let $X$ be a toric variety and $f\colon{\mathcal{C}}\rightarrow X$ a family of parameterized curves defined over the field $K$. Assume that the base $B$ of the family is quasi-projective, and consider the set of tropicalizations
$\Sigma:=\left\{\left[\mathrm{trop}(f_{\eta},C_{\eta})\right]\;:\;\eta\in B(K)\right\}\subseteq M^{\mathrm{trop}}_{g,n,\nabla}.$
@@
    If $M_{[\Theta]}$ is weightless and $3$-valent and $\dim(\overline{\Sigma}\cap M_{[\Theta]})=\dim M_{[\Theta]}$, then $\overline{M}_{[\Theta]}\subseteq\overline{\Sigma}$;
  3. (3)
In [CHT25] we apply this corollary to prove the irreducibility of Severi varieties parametrizing integral curves of a given geometric genus on certain polarized toric surfaces. We then deduce the irreducibility of the classical Hurwitz spaces in any characteristic. More precisely, in loc. cit., we apply the corollary to the universal family over an irreducible component $B$ of the Severi variety. By imposing $\dim(B)$ point constraints, whose tropicalizations are in tropically general position, one can easily show that there exists a weightless and $3$-valent combinatorial type $\Theta$ such that $\dim(\overline{\Sigma}\cap M_{[\Theta]})=\dim M_{[\Theta]}$. On the other hand, one can show that there exists a weightless and $3$-valent type $\Theta_{0}$ such that $\overline{M}_{[\Theta_{0}]}\subseteq\overline{\Sigma}_{0}$, where $\Sigma_{0}$ corresponds to a unique irreducible component $B_{0}$ of the Severi variety. Thus, Assertion (2) of the corollary, reduces the irreducibility problem to a purely combinatorial question reminiscent of establishing connectivity of the moduli space $M^{\mathrm{trop}}_{g,n,\nabla}$ in codimension one. Namely, is there a sequence of (almost) $3$-valent strata adjacent to each other starting at $M_{[\Theta]}$ and ending at $M_{[\Theta_{0}]}$?
Another application of the results of the current paper will be a computation of the number of moduli of curves in (irreducible) Severi varieties. More precisely, normalizing the universal family over a Severi variety $V$ induces a map from $V$ to the moduli space ${\mathcal{M}}_{g}$ of smooth genus $g$ curves. Similarly, on the tropical side, we obtain a natural map from the moduli space $M_{g,\nabla}^{\mathrm{trop}}$ to the moduli space $M_{g}^{\mathrm{trop}}$ of stable tropical curves of genus $g$. Our main result, together with the results of [ACP15] for $M_{g}^{\mathrm{trop}}$, allows to relate the tropical and the algebro-geometric side of this picture. In a forthcoming paper [CHT26] we use it to establish the dimension of the image of Severi varieties in the moduli space ${\mathcal{M}}_{g}$.
Acknowledgements: We thank an anonymous referee for many thoughtful comments.
@@ 2. Notation and terminology
Throughout this paper, we work over a valued field $(K,\nu)$, which is the algebraic closure of a complete discretely valued field $F$ with algebraically closed residue field. The ring of integers of $K$ is denoted by $K^{0}$, the maximal ideal by $K^{00}$, and the residue field by $\widetilde{K}$; similarly for $F$.
Let $N\simeq{\mathbb{Z}}^{n}$ be a lattice, and $N_{\mathbb{R}}:=N\otimes_{\mathbb{Z}}{\mathbb{R}}$ the corresponding vector space. By a polyhedron in $N_{\mathbb{R}}$ we mean the intersection of finitely many affine half-spaces. A polyhedron is said to have rational slopes if the presentation above can be chosen such that all half-spaces are given by integral normal vectors. For a subset $P\subset N_{\mathbb{R}}$, we denote by ${\rm Lin}(P)$ the linear subspace acting simply transitively on the affine hull of $P$. If $P$ is a polyhedron with rational slopes, then ${\rm Lin}(P)$ admits a natural integral structure induced from $N_{\mathbb{R}}$, i.e., the lattice $N\cap{\rm Lin}(P)$ of integral vectors. By an integral structure on $P$, we mean the integral structure on ${\rm Lin}(P)$.
@@ 2.1. Polyhedral complexes
An (abstract) polyhedral complex with integral structure is a connected topological space $\Lambda$ together with a finite set $\mathcal{F}(\Lambda)$ of closed subsets, called faces of $\Lambda$, and a map $\mu_{W}\colon W\rightarrow N_{W}\otimes{\mathbb{R}}$ for every $W\in\mathcal{F}(\Lambda)$ such that
@@
    $N_{W}$ is a lattice of finite rank for any face $W\in\mathcal{F}(\Lambda)$;
@@
    The map $\mu_{W}$ is a homeomorphism from $W$ to a polyhedron with rational slopes of dimension ${\rm rank}(N_{W})$. By abuse of language, we will address $W$ itself as a polyhedron and speak about its faces, interior, dimension, integral structure, etc.;
@@
    If $W\in\mathcal{F}(\Lambda)$ and $W^{\prime}\in\mathcal{F}(\Lambda)$ is a face of $W$, then the integral structure on $W^{\prime}$ is the restriction of the integral structure on $W$, i.e., the linear part of $\mu_{W}\circ\mu_{W^{\prime}}^{-1}$ takes $N_{W^{\prime}}$ to a saturated sublattice in $N_{W}$.
We say that $\Lambda$ is rational if each $\mu_{W}(W)$ has rational vertices in $N_{W}\otimes{\mathbb{R}}$. All polyhedral complexes considered in this paper will be assumed to have integral structures, and we will call them polyhedral complexes for short.
@@
A map $\beta\colon\Lambda\rightarrow\Lambda^{\prime}$ is called piecewise integral affine if for each face $W\in\mathcal{F}(\Lambda)$ there is a face $W^{\prime}\in\mathcal{F}(\Lambda^{\prime})$ such that $\beta(W)\subseteq W^{\prime}$ and $\beta|_{W}\colon W\to W^{\prime}$ is integral affine.
For a face $W\in\mathcal{F}(\Lambda)$, let $\{W_{i}\}_{i}$ be the set of faces in $\mathcal{F}(\Lambda)$ containing $W$ as a codimension-one face. Then, for each $i$, the lattice $N_{W}$ is naturally identified with a saturated sublattice of $N_{W_{i}}$ of corank one. Let $\vec{e}_{i}+N_{W}$ be the generator of $N_{W_{i}}/N_{W}$ for which $\mu_{W_{i}}(W_{i})\subseteq\mu_{W_{i}}(W)+\mathbb{R}_{\geq 0}\vec{e}_{i}$, and denote by ${\rm Star}(W)$ the collection $\{\vec{e}_{i}+N_{W}\}$. We will often consider a small open neighborhood of $W^{\circ}$ in $W^{\circ}\cup\left(\bigcup_{i}W_{i}^{\circ}\right)$. To distinguish this set from ${\rm Star}(W)$, we denote it by ${\rm Star}(W^{\circ})$.
@@
Let $\Lambda$ be a polyhedral complex and $W\in\mathcal{F}(\Lambda)$ a face. A piecewise integral affine map $\beta\colon\Lambda\rightarrow{\mathbb{R}}^{n}$ is called quasi-harmonic at $W$ if
$\sum_{\vec{e}+N_{W}\in{\rm Star}(W)}a_{\vec{e}}\cdot\frac{\partial\beta}{\partial\vec{e}}\in{\rm Lin}(\beta(W))$
for some positive integers $a_{\vec{e}}$. It is called harmonic at $W$ if one can choose all $a_{\vec{e}}=1$. Finally, the map $\beta$ is called (quasi-)harmonic if it is (quasi-)harmonic at all codimension-one faces.
@@
Plainly, being (quasi-)harmonic at $W$ depends only on the restriction of the map $\beta$ to ${\rm Star}(W^{\circ})$.
@@ 2.2. Tropical curves
We next recall some conventions concerning tropical curves and their moduli from [CHT23, CHT22], and extend the notion of one-parameter families of parameterized tropical curves therein to arbitrary dimension of the base.
@@ 2.2.1. Abstract and parameterized tropical curves
A tropical curve $\Gamma$ consists of an underlying finite graph $\mathbb{G}$}, a length function $\ell\colon E(\mathbb{G})\rightarrow\mathbb{R}_{>0}$ and a weight (or genus) function $g\colon V(\mathbb{G})\rightarrow\mathbb{Z}_{\geq 0}$. Here $V({\mathbb{G}})$ denotes the set of vertices of $\mathbb{G}$}, and $E({\mathbb{G}})$ the set of edges. The graph $\mathbb{G}$} is furthermore equipped with an ordered set of half-edges, which we call legs. The set of legs is denoted by $L({\mathbb{G}})$, and we define $\overline{E}(\mathbb{G}):=E({\mathbb{G}})\cup L({\mathbb{G}})$. We extend $\ell$ to the legs $l\in L({\mathbb{G}})$ by setting $\ell(l):=\infty$.
We identify tropical curves with their geometric realization as polyhedral complexes, without necessarily distinguishing the two viewpoints. That is, we may divide the loops at the middle point, and then identify the edges of $\mathbb{G}$} with bounded closed intervals of the corresponding lengths, and the legs of $\mathbb{G}$} with semi-bounded closed intervals.
For $e\in\overline{E}({\mathbb{G}})$, we write $\vec{e}$ to indicate a choice of orientation on $e$. If $e$ is a leg, then it will always be oriented away from the vertex, while bounded edges will be considered with both possible orientations. In particular, the elements in ${\rm Star}(v)$ are in one-to-one correspondence with the collection of oriented edges and legs having $v$ as their tail. The number of elements in ${\rm Star}(v)$ is called the valence of $v$. The genus of $\Gamma$ is defined to be $g(\Gamma)=g({\mathbb{G}}):=1-\chi(\mathbb{G})+\sum_{v\in V(\mathbb{G})}g(v)$, where $\chi({\mathbb{G}}):=b_{0}({\mathbb{G}})-b_{1}({\mathbb{G}})$ is the Euler characteristic of $\mathbb{G}$}.
A tropical curve $\Gamma$ is called weightless if the weight function is identically zero and stable if $|{\rm Star}(v)|+2g(v)\geq 3$ for every vertex $v\in V({\mathbb{G}})$.
To define parameterized tropical curves, we fix a lattice $N$ and set $N_{\mathbb{R}}:=N\otimes{\mathbb{R}}$. Then a parameterized tropical curve is, by definition, a harmonic piecewise integral affine map $h\colon\Gamma\rightarrow N_{\mathbb{R}}$ from a tropical curve $\Gamma$ (considered as a polyhedral complex) to $N_{\mathbb{R}}$. The combinatorial type $\Theta$ of a parameterized tropical curve is defined to be the underlying weighted graph $\mathbb{G}$ together with the collection of slopes $\frac{\partial h}{\partial\vec{e}}$ for $e\in\overline{E}({\mathbb{G}})$. The extended degree $\overline{\nabla}$ of a parameterized tropical curve consists of the sequence of slopes $\left(\frac{\partial h}{\partial\vec{l}}\right)_{l\in L({\mathbb{G}})}$, and the degree $\nabla$ of the subsequence $\left(\frac{\partial h}{\partial\vec{l}_{i}}\right)$ of non-zero slopes.